OpenStudy (anonymous):

MEDAL AND FAN The equation below shows the perimeter f(x) of several rectangular side tables having width x inches: f(x) = 24 + 2x The perimeter of each table is less than 50 inches. Which of the following best describes the domain of f(x)? A)All integers greater than 0 and less than 13 B) All integers greater than 0 and less than 50 C) All real numbers greater than 0 and less than 13 D) All real numbers greater than 0 and less than 50

OpenStudy (anonymous):

f(x) <=50 f(x) = 24 + 2x 50 = 24 + 2x -24 -24 26=2x 26/2 2x/2 13=x

OpenStudy (anonymous):

That is what I have so far @misty1212 could you help me please?

OpenStudy (anonymous):

I know its not B or D

OpenStudy (anonymous):

Im thinking A

OpenStudy (misty1212):

i think you did it only the inequality became an equal sign

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

So it is A? I got it right?

OpenStudy (misty1212):

greater than zero, less than 13 right?

OpenStudy (anonymous):

Yes integers correct?

OpenStudy (misty1212):

no i don't think they have to be integers

OpenStudy (misty1212):

then could be fractions too

OpenStudy (misty1212):

pick C C) All real numbers greater than 0 and less than 13

OpenStudy (anonymous):

I would think it would be A or C integers or Real numbers but I chose integers based on the theory it could be anything

OpenStudy (anonymous):

Okies :3 Thankies!

OpenStudy (misty1212):

one side could be \(10.5\) or some other number that is not an integer

OpenStudy (misty1212):

welkies (?)

OpenStudy (anonymous):

Thats true.

OpenStudy (anonymous):

my view on it is A doesnt say it is only integers but says it accepts all integers between those numbers

OpenStudy (anonymous):

But I agree with you